Showing all work, determine whether the lines given by the equations below are parallel, perpendicular, or neither:
y = 2/3x - 6
6x + 4y = 5

Answer by rfer(16322)

y=2/3x-6
----
6x+4y=+5
4y=-6x+5
y=-3/2x+5/4
--------------
Perpendicular